The line through A(x,5) and B(4,3) is perpendicular to a line whose slope is 3/5. Find the value of x.
Answer by checkley71(8403)

ALINE PERPENDICULAR TO A LINE WITH A SLOPE OF 3/5 IS A SLOPE OF -5/3
THUS THE SLOPE PASSING THROUGH THE POINTS (X,5)(4,3) IS -5/3
(3-5)/(4-X)=-5/3
-2/(4-X)=-5/3 CROSS MULTIPLYING WE GET
-6=-5(4-X)
-6=-20+5X
5X=-6+20
5X=14
X=14/5
PROOF
(3-5)/(4-14/5)=-5/3
-2/(20-14)/5=-5/3
-2/(6/5)=-5/3
-2*5/6=-5/3
-10/6=-5/3
-5/3=-5/3